OpenStudy (anonymous):

Use the function to evaluate the indicated expressions and simplify. f (x) = 7x - 1 f (2x) = _____ 2f (x) = _____

OpenStudy (anonymous):

If f(x) = 7x - 1 Then f(2x) would be 7(2x) - 1 If f(x) is 7x - 1 Then 2 times f(x) would be 2(7x -1) Just simplify those two expressions and you have your answer.

OpenStudy (anonymous):

so they are 14x-7 and 14x-7

OpenStudy (anonymous):

?

OpenStudy (anonymous):

14x-1 and 14x-7

OpenStudy (anonymous):

i mean 14x-2

OpenStudy (anonymous):

why would it be 14x-1

OpenStudy (anonymous):

Sorry 14x-2

OpenStudy (anonymous):

Yeap 14x-1 and 14x-2

OpenStudy (anonymous):

that doesnt make sense if you distribute

OpenStudy (anonymous):

7(2x) - 1 = 14x - 1 2(7x - 1) = 2(7)x - 2(1) = 14x - 2

OpenStudy (anonymous):

actually you plug in the f(2x) into the x part of the equation so: f(2x)=7x-1=7(2x)-1=14x-1

OpenStudy (anonymous):

yup I see now some others already covered this